OpenStudy (anonymous):

Help? Please? How would I simplify x^-6/x^-3 also how would I write that with positive exponents?

OpenStudy (anonymous):

try to do it. I am watching

OpenStudy (anonymous):

I'm not sure how though

OpenStudy (anonymous):

nothing to switch from numerator, but the denominator can. I got something like x^-6*x^3. can you continue?

OpenStudy (anonymous):

ok, go another way, to have positive,

OpenStudy (anonymous):

I switch x^-6 from numerator into denominator, so I change the sign of exponential. is it ok?

OpenStudy (anonymous):

I get \[\frac{ 1 }{ x^6x^-3 } = \frac{ 1 }{ x^3 }\]

OpenStudy (anonymous):

I believe it is the right answer which satisfy your requirement. (simplify and get positive exponential)

OpenStudy (anonymous):

Thanks so much

OpenStudy (anonymous):

yw